nominal gdp in year 1 is 500. if gdp deflator doubles in year 8 while real output increases by 60 percent, what's nominal output in year 8


1
Expert's answer
2021-06-22T15:37:03-0400

Assuming 1 year as a base year 

 GDP Deflator= (Nominal GDP)"\div" (Real GDP)x100

   "x=(500\times× 100\times× 100)\div 60"

  "x=83,333.33"

As mentioned above, if GDP deflator doubles in year 8 while real output increased by 60%

then GDP deflator in year"8= 83,333.33\times 2"

                  "= 166,666.66"

therefore Nominal GDP in year 8= (GDP deflator in year 8"\times" real GDP)"\div" 100

                           "=(166,666.66\times0.6)\div 100"

                           "=999.99"

Nominal output= (Nominal GDP"\div" Real GDP)(100)

             "= 999.99\div 0.6(100)"

             "= 166,665"

The nominal output in 8 year will be = 166,665